# Discovering modular solutions that generalize compositionally

### v1

@@ B.4 Compositional goals
Over-parameterization in the compositional goal environment.
Similarly, we investigated the effect of varying the hidden dimension and number of modules in the hypernetwork models on the OOD accuracy achieved in the compositional goals environment. Different from the compositional preference environment, we have a discrete set of goals in this case. While in the discrete multi-task teacher-student setting OOD performance was sensitive to over-parameterization, Figure A10 shows relatively stable OOD accuracy across varying hidden and module dimension.
@@ C.1.1 Data generation
For all experiments in the multi-task teacher student, we fix the activation function $\psi$ to the ReLU nonlinearity, the task latent variable dimension to $M=6$, input dimension to $n=16$, hidden dimension of the teacher to $h=16$, and output dimension to $o=4$. The teacher parameters $\Theta$ and $\bm{a}$} are generated by sampling the entries i.i.d. from the centered truncated normal distribution, with standard deviation resp. $\frac{1}{\sqrt{M}}$ and $\frac{1}{\sqrt{h}}$. We define the distribution over inputs $\bm{x}$} to be the uniform distribution with 0 mean and standard deviation $1$ on $\mathbb{R}^{n}$. Finally, we specify the distribution over task latent variable $\bm{z}$}. We distinguish 2 settings:
@@
Here, we consider tasks where modules are sparsely, linearly combined. A task distribution is specified by a set of masks, that are binary vectors of $\mathbb{R}^{M}$. Given a mask, we sample a task $\bm{z}$} as follows. We first sample an $M$-dimensional random variable following the exponential distribution. Then, we zero-out entries corresponding to the mask being 0. We then normalize the vector such that the sum equals 1. This procedure simulates the uniform sampling from the simplex spanning the directions in which the mask is non-zero. Finally, we add the mask to the vector and rescale the outcome by $0.5$. This ensures two tasks generated by distinct masks do not have intersecting support (but intersecting span). See Algorithm 1 for the pseudocode.
@@ C.1.5 Hyperparameters
For all experiments in the multi-task teacher student, we set $B_{\text{outer}}=64$, $B_{\text{inner}}=256$, $N_{\text{outer}}=60000$, $N_{\text{inner}}=300$. We optimize the inner loop using the Adam optimizer with $\eta_{\text{inner}}=0.003$, and the outer loop using the AdamW optimizer with various weight decay strengths and an initial learning rate $\eta_{\text{outer}}=0.001$ annealed using cosine annealing down to $10^{-6}$ at the end of training. We observed that this set of hyperparameters gave enough time for the student to practically reach 0 loss, which is the setting studied in the theory.
@@ C.2.1 Data generation
We generate data by first initializing the teacher parameters once. The teacher network is a linear hypernetwork for which the parameterization and initialization scheme are described in Section D. For all our experiments, unless specified otherwise, we fix the input dimension to $n=16$, hidden dimension of the teacher to $h=32$, and output dimension to $o=8$. We define the distribution over inputs $\bm{x}$} to be the uniform distribution with 0 mean and standard deviation $1$ on $\mathbb{R}^{n}$.
Targets are generated by passing the input $\bm{x}$} through the generated network. To make sure the distribution of the outputs is not too dissimilar across tasks, we normalize each output neurons to have unit variance and zero mean by numerically estimating the first and second moment of the distribution over random inputs. The tasks are sampled following the same procedure as in the discrete task distribution description in Section C.1.1. That is, a finite set of binary vectors of size $M$ fully specifies the task distribution.
@@ C.2.3 Experiments
We perform this experiment for different values of $K$. Given $K$, we compute the set of binary masks with up to $K$ non-0 entries, and randomly select three-quarters of them and use the obtained list for generating training tasks. We make sure that the obtained list has compositional support, that is, all modules appear at least once. The remaining quarter is used as the OOD task for evaluation. For non-compositional support, we again enumerate all binary masks with up to $K$ non-0 entries, but split it into 2 sets such that for training, only masks where the last module is inactive are used, while for the evaluation, masks which always have the last module activated, are used to generate the tasks.
@@ C.3.3 Experiments
For compositional support, we compute the set of binary masks with up to 33 non-0 entries, and randomly select three-quarter of them and use the obtained list for generating training tasks. We make sure that the obtained list contain a compositional support. The remaining quarter is used as the OOD task for evaluation. For non-compositional support, we enumerate all binary masks with up to 33 non-0 entries, and split it into 2 sets following the value of the last entry of the mask. For training, we use the set of mask where the last module is inactive, while for the evaluation, we use the set of mask which always has the last module activated, to generate the tasks. Again, these settings were chosen such that the number of tasks for the compositional and non-compositional experiment are comparable.
@@ C.4.1 Data generation
For the compositional goals environment we consider mazes of size $11\times 11$ with the same movement dynamics as described above in the compositional preference environment. In addition to moving up, down, left and right, the agent can choose one of the ‘object interaction’ actions. Each environment has 5 randomly placed objects, one of which is the target object. The agent is placed randomly and has to reach the target object followed by performing the target interaction to obtain a reward. Each task is defined by a goal vector that specifies which of the 5 maze configurations is used, which of the 5 object types is assigned as the target object, which of the 2 possible target interactions is the correct one and finally in which of the 4 quadrants of the maze the target object will be located (c.f. Figure 4D.)
@@ D.1.2 Initialization
We initialize $\phi_{0}$ such that parameters corresponding to weights are initialized following the truncated normal distribution with mean 0 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer. Parameters corresponding to biases are initialized at 0.
@@ D.2.2 Initialization
We initialize $\theta$ and $\phi$ such that parameters corresponding to weights are initialized following the truncated normal distribution with mean 0 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer. Parameters corresponding to biases are initialized at 0.
@@ D.3.2 Initialization
The hypernetwork weights are initialized following the truncated normal distribution with mean 0 and standard deviation $\frac{1}{\sqrt{M}}$ where $M$ is the embedding vector dimension.
The meta-parameters corresponding to the initialization values of the embeddings are initialized following a uniform and unit variance distribution, and those corresponding to biases are initialized to 0.
@@ D.4.2 Initialization
The hypernetwork parameters are initialized such that all biases are set to 0 and MLP layer weights are initialized following the truncated normal distribution with mean 0 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer.
The meta-parameter corresponding to the initialization value of the embeddings are initialized following a uniform and unit variance distribution, and those corresponding to biases are initialized to 0.

### v2

@@ B.4 Compositional goals
Overparameterization in the compositional goal environment.
Similarly, we investigated the effect of varying the hidden dimension and number of modules in the hypernetwork models on the OOD accuracy achieved in the compositional goals environment. Different from the compositional preference environment, we have a discrete set of goals in this case. While in the discrete multi-task teacher-student setting OOD performance was sensitive to overparameterization, Figure A10 shows relatively stable OOD accuracy across varying hidden and module dimension.
@@ C.1.1 Data generation
For all experiments in the multi-task teacher student, we fix the activation function $\psi$ to the ReLU nonlinearity, the task latent variable dimension to $M=6$, input dimension to $n=16$, hidden dimension of the teacher to $h=16$, and output dimension to $o=4$. The teacher parameters $\Theta$ and ${\bm{a}}$ are generated by sampling the entries i.i.d. from the centered truncated normal distribution, with standard deviation resp. $\frac{1}{\sqrt{M}}$ and $\frac{1}{\sqrt{h}}$. We define the distribution over inputs ${\bm{x}}$ to be the uniform distribution with 00 mean and standard deviation $1$ on $\mathbb{R}^{n}$. Finally, we specify the distribution over task latent variable ${\bm{z}}$. We distinguish 2 settings:
@@
Here, we consider tasks where modules are sparsely, linearly combined. A task distribution is specified by a set of masks, that are binary vectors of $\mathbb{R}^{M}$. Given a mask, we sample a task ${\bm{z}}$ as follows. We first sample an $M$-dimensional random variable following the exponential distribution. Then, we zero-out entries corresponding to the mask being 00. We then normalize the vector such that the sum equals 1. This procedure simulates the uniform sampling from the simplex spanning the directions in which the mask is non-zero. Finally, we add the mask to the vector and rescale the outcome by $0.5$. This ensures two tasks generated by distinct masks do not have intersecting support (but intersecting span). See Algorithm 1 for the pseudocode.
@@ C.1.5 Hyperparameters
For all experiments in the multi-task teacher student, we set $B_{\text{outer}}=64$, $B_{\text{inner}}=256$, $N_{\text{outer}}=60000$, $N_{\text{inner}}=300$. We optimize the inner loop using the Adam optimizer with $\eta_{\text{inner}}=0.003$, and the outer loop using the AdamW optimizer with various weight decay strengths and an initial learning rate $\eta_{\text{outer}}=0.001$ annealed using cosine annealing down to $10^{-6}$ at the end of training. We observed that this set of hyperparameters gave enough time for the student to practically reach 00 loss, which is the setting studied in the theory.
@@ C.2.1 Data generation
We generate data by first initializing the teacher parameters once. The teacher network is a linear hypernetwork for which the parameterization and initialization scheme are described in Section D. For all our experiments, unless specified otherwise, we fix the input dimension to $n=16$, hidden dimension of the teacher to $h=32$, and output dimension to $o=8$. We define the distribution over inputs ${\bm{x}}$ to be the multivariate uniform distribution over the hypercube $[-1,1]^{n}$.
Targets are generated by passing the input ${\bm{x}}$ through the generated network. To make sure the distribution of the outputs is not too dissimilar across tasks, we normalize each output neuron to have unit variance and zero mean by numerically estimating the first and second moment of the distribution over random inputs. The tasks are sampled following the same procedure as in the discrete task distribution description in Section C.1.1. That is, a finite set of binary vectors of size $M$ fully specifies the task distribution.
@@ C.2.3 Experiments
We perform this experiment for different values of $K$. Given $K$, we compute the set of binary masks with up to $K$ non-00 entries, and randomly select three-quarters of them and use the obtained list for generating training tasks. We make sure that the obtained list has compositional support, that is, all modules appear at least once. The remaining quarter is used as the OOD task for evaluation. For non-compositional support, we again enumerate all binary masks with up to $K$ non-00 entries, but split it into 2 sets such that for training, only masks where the last module is inactive are used, while for the evaluation, masks which always have the last module activated, are used to generate the tasks.
@@ C.3.3 Experiments
For compositional support, we compute the set of binary masks with up to $3$ non-00 entries, and randomly select three-quarter of them and use the obtained list for generating training tasks. We make sure that the obtained list contain a compositional support. The remaining quarter is used as the OOD task for evaluation. For non-compositional support, we enumerate all binary masks with up to $3$ non-00 entries, and split it into 2 sets following the value of the last entry of the mask. For training, we use the set of mask where the last module is inactive, while for the evaluation, we use the set of mask which always has the last module activated, to generate the tasks. Again, these settings were chosen such that the number of tasks for the compositional and non-compositional experiment are comparable.
@@ C.4.1 Data generation
For the compositional goals environment we consider mazes of size $11\times 1$ 11 with the same movement dynamics as described above in the compositional preference environment. In addition to moving up, down, left and right, the agent can choose one of the ‘object interaction’ actions. Each environment has 5 randomly placed objects, one of which is the target object. The agent is placed randomly and has to reach the target object followed by performing the target interaction to obtain a reward. Each task is defined by a goal vector that specifies which of the 5 maze configurations is used, which of the 5 object types is assigned as the target object, which of the 2 possible target interactions is the correct one and finally in which of the 4 quadrants of the maze the target object will be located (c.f. Figure 4D.)
@@ D.1.2 Initialization
We initialize $\phi_{0}$ such that parameters corresponding to weights are initialized following the truncated normal distribution with mean 00 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer. Parameters corresponding to biases are initialized at 00.
@@ D.2.2 Initialization
We initialize $\theta$ and $\phi$ such that parameters corresponding to weights are initialized following the truncated normal distribution with mean 00 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer. Parameters corresponding to biases are initialized at 00.
@@ D.3.2 Initialization
The hypernetwork weights are initialized following the truncated normal distribution with mean 00 and standard deviation $\frac{1}{\sqrt{M}}$ where $M$ is the embedding vector dimension.
The meta-parameters corresponding to the initialization values of the embeddings are initialized following a uniform and unit variance distribution, and those corresponding to biases are initialized to 00.
@@ D.4.2 Initialization
The hypernetwork parameters are initialized such that all biases are set to 00 and MLP layer weights are initialized following the truncated normal distribution with mean 00 and standard deviation $\frac{1}{\sqrt{H}}$ where $H$ is the size of the input to the layer.
The meta-parameter corresponding to the initialization value of the embeddings are initialized following a uniform and unit variance distribution, and those corresponding to biases are initialized to 00.